# Out-of-bag prediction balls for random forests in metric spaces

### v1

@@ 3.3.4 Probability distributions on the real line
The same idea characterizes the Fréchet mean of $\mathrm{P}\in\mathcal{W}_{2}(\mathbb{R})$ as the unique measure $\mathrm{P}_{\oplus}\in\mathcal{W}_{2}(\mathbb{R})$ whose quantile function satisfies $F_{\mathrm{P}_{\oplus}}^{-1}=\mathsf{E}(F_{\mathrm{P}}^{-1})$ (see, e.g., Theorem 3.2.11 of Panaretos and Zemel,, 2020), thus ensuring (c.2)(c.2.1).
The definition of regression models with response on $W_{2}(\mathbb{R})$ is relatively straightforward. For a given probability distribution $\mathrm{Y}\in\mathcal{W}_{2}(\mathbb{R})$, consider a slight modification of the regression model in Petersen and Müller, (2019, Section 6.2):
@@
for some $\mu\in\mathcal{W}_{2}(\mathbb{R})$ and some finite real functions $f$ and $g$, under the restriction that $\sigma_{0}+g(x)>0$ for all $x$ in the support of $X$, so that $m(x)(\cdot)$ is a valid quantile function in $L^{2}[0,1]$. In order to generate the response, consider $F_{\mathrm{Y}}^{-1}(\cdot)=\gamma+f(X)+(\sigma+g(X))F_{\mu}^{-1}(\cdot)$, with $\gamma$ and $\sigma$ random variables independent of $X$ and following distributions of mean $\gamma_{0}$ and $\sigma_{0}$, respectively. See Serrano and García-Portugués, (2025) for a detailed numerical example with this regression model. Thus, this general model satisfies (c.2)(c.2.2), and the conditional Fréchet mean is $m_{0}(x)(\cdot)$. The population OOB balls in this setting are formed by the distributions whose quantile functions are within a constant $L_{2}$-distance on $[0,1]$ equal to $R_{1-\alpha}$. If we assume, for simplicity, that $F_{\mu}^{-1}(\cdot)$ is the quantile of a distribution with mean zero and unit variance, then $R_{1-\alpha}$ is the $(1-\alpha)$-quantile of $((\gamma-\gamma_{0})^{2}+(\sigma-\sigma_{0})^{2})^{1/2}$.
@@ 5 Numerical experiments
In this section, we assess the performance of OOB balls through simulations with responses in the Euclidean space $\mathbb{R}^{q}$, on the unit sphere $\mathbb{S}^{2}$ and unit hyperboloid $\mathbb{H}^{2}$, and in the space of SPD matrices $\mathcal{S}_{2}^{+}$. In $\mathbb{R}^{q}$, OOB balls are compared with SC balls for metric data (Lugosi and Matabuena,, 2024) in terms of coverage (Types I–IV), radius, and computational time.
@@
Each tree was trained with a size $n$ subsample generated with nonparametric bootstrap with replacement. We used $B=200$ trees, so that the number of trees in each OOB prediction is closer to the default parameter $B=100$ in scikit-learn (Pedregosa et al.,, 2011) and works related to RFs in metric spaces (Bulté and Sørensen,, 2024; Qiu et al.,, 2024), without significantly increasing the computational cost of fitting and predicting. We investigated the effect of increasing $B$ on Type I coverage results, for OOB and SC balls, and found no significant improvement in coverage when the number of trees exceeded $B=200$.
The CART criterion was selected as impurity measure, and the split method used was the $2$-means algorithm. The minimal split size, which determines the minimal number of samples allowed in the leaf nodes, was tuned from $\{1,5,10\}$ using grid-search $5$-fold cross-validation with the MSE as error measure. Using the same cross-validation procedure, the number of random candidate variables to consider at each split was tuned from $\{1,2,\ldots,p\}$. This parameter is denoted by mtry in the pyfréchet package and in the most popular R programming language implementations (Liaw and Wiener,, 2002; Wright and Ziegler,, 2017). Three different values were considered for the significance level $\alpha$: $0.01,0.05$, and $0.10$.
@@ 5.2 Sphere
In the subsequent simulation scenarios, the focus is on the performance of OOB balls (radius and coverage types). We used RFWLCFR to train RFs using the implementation from Bulté and Sørensen, (2024) in the pyfréchet package for Python. For the unit sphere $\mathbb{S}^{2}$, we considered the regression function $\boldsymbol{m}_{0}$ (great circle) given in (13), with $\boldsymbol{\mu}=(1/\sqrt{2},1/\sqrt{2})^{\top}$. The predictor variable $\Theta$ was generated as $\Theta=\operatorname{arctan2}(X_{2},X_{1})$, where $\boldsymbol{X}=(X_{1},X_{2})\sim\mathrm{vMF}((1,0)^{\top},1)$. For the response, we considered $\boldsymbol{Y}\mid\Theta=\theta\sim\mathrm{vMF}(\boldsymbol{m}_{0}(\theta),\kappa)$ for $\kappa=50$, $200$. The data generation process is illustrated in Figure 2 for different values of $\Theta$.
@@ 5.4 Symmetric positive definite matrices
We consider the space $\mathcal{S}_{q}^{+}$ of SPD $q\times q$ matrices, and study the impact of metric selection on the performance of OOB balls. Specifically, we compared the coverage and radius of OOB balls in $\mathcal{S}_{2}^{+}$ using the three metrics. The regression model is constructed based on the idea of a convex interpolation between SPD matrices introduced in Section 3.3.3. In particular, we considered the interpolation function
@@ 7 Out-of-bag balls to model sunspot dynamics
Due to its plasma composition, the Sun’s equatorial regions rotate faster than the polar regions (differential rotation). This phenomenon distorts the magnetic field, stretching the Sun’s initial poloidal magnetic field into toroidal configurations, as plasma drags field lines along their rotational trajectories. Each solar cycle lasts around $11$ years and culminates when these twisted fields undergo a global polarity reversal (see, e.g., Babcock,, 1961). The entanglement of field lines creates localized regions of intensified magnetic pressure that suppress convective heat transfer, manifesting as cooler, darker sunspots. Sunspots have a variable lifetime and serve as direct observational tracers of solar activity, which is linked to space weather events that impact spacecraft, astronauts, and Earth’s infrastructure.
We analyzed data based on a merged catalog combining the Greenwich Photoheliographic Results (GPR) and the Debrecen Photoheliographic Data (DPD) sunspot records (Baranyi et al.,, 2016; Győri et al.,, 2016). The GPR spans the years 1872–1976, while the DPD extends the record from 1974 onward, covering 14 solar cycles to date. The R package rotasym (García-Portugués et al.,, 2025) provides the data sets sunspots_births and sunspots_deaths, containing, respectively, the location on the Sun of the first and last record of each sunspot group (referred to as “birth” and “death” henceforth). We merged these data sets and excluded the first ($11$th) and last ($24$th) cycles due to the low reliability of their observations. Sunspots that were born and died on the same day were also excluded.
@@
For each solar cycle, we predicted death locations from birth locations using RFWLCFR, treating both response and predictors as points on the unit sphere $\mathbb{S}^{2}$ with geodesic distance $d_{\mathbb{S}^{2}}$. Although Qiu et al., (2024) consider only real predictors, their method can be applied straightforwardly when the predictors belong to any metric space. The data from each cycle were split into training ($75\%$) and test ($25\%$) sets, and the RFs were tuned with the same parameters as in Section 5, using the implementation in pyfréchet. Although the RF captured the latitude-dependent directional pattern (Figure 6(b)), the predicted longitudinal displacements are systematically smaller than observed in the training data. This shrinkage effect likely arises from the model’s averaging of opposing trajectories present at all latitudes, a consequence of the chaotic motion of sunspots.
@@
The predictors remained on $\mathbb{S}^{2}$ with the geodesic distance $d_{\mathbb{S}^{2}}$, while the responses were modeled on $S_{a,c}^{2}$, equipped with its intrinsic geodesic distance. Karney, (2013) provides an algorithm to compute $d_{S_{a,c}^{2}}$ which can be made arbitrarily accurate and always converges. Although $d_{S_{a,c}^{2}}$ defines a Riemannian manifold structure on $S_{a,c}^{2}$, to simplify the implementation, we employed Fréchet medoids instead of Fréchet means in the CART criterion (4) and in the aggregation of results (i.e., we minimized (3) over the sample elements). On the sphere ($a=c=1$), medoids were also used for the sake of homogeneity in the comparison.
@@
Table 6: For each specification of $a$ (and $c=1$), the errors, areas, and coverages were calculated on the test points of cycles 21212323, and the reported values are weighted means across the three cycles, with weights proportional to the number of test points on each cycle. The areas correspond to OOB balls centered on the predicted value of the test points. The relative MSE difference $\Delta_{\mathrm{MSE}}$ and the relative mean area difference $\Delta_{\mathrm{area}}$ (both in $\%$) were calculated with respect to the balls on the sphere. The reported coverage corresponds to Type II (in %) for $\alpha=0.10$. The $p$-value of the one-sided paired $t$-test with alternative hypothesis $H_{1}:\mathrm{MSE}_{\mathbb{S}^{2}}<\mathrm{MSE}_{S_{a,1}}$ is $p_{\mathrm{MSE}}$. For $p_{\mathrm{area}}$, we considered a one-sided paired $t$-test with alternative $H_{1}:\mathrm{area}_{\mathbb{S}^{2}}>\mathrm{area}_{S_{a,1}}$, to test the equality of mean areas. False discovery rate correction Benjamini and Yekutieli, (2001) is applied to the $p$-values at level $\alpha=0.01$.
@@ 8 Discussion
In this work, a novel approach is introduced to quantify the uncertainty in RF predictions for metric data. The nature of a RF organically provides the OOB balls, with a main advantage over existing SC methods: the entire sample is used to train the RFs that generate the OOB balls, and to estimate the ball radius. Thus, OOB balls need half the sample size that standard SC methods require to achieve the same prediction and inference accuracy, as corroborated in simulations. Although OOB balls reduce the effective number of trees by a factor of $e^{-1}$, this can be compensated by increasing the total number of trees.
A limitation of conformal prediction is the elevated cost of in-sample inference. Suppose one wants to analyze the confidence regions at the same points where the RF was trained. Although rank-one-out procedures (Lei et al.,, 2018) reduce significantly the cost of in-sample SC inference, this method inherits the inefficiency of SC prediction due to halving the sample size, and the computational cost is doubled since one RF is trained on each half sample. A possible competitive extension is to apply OOB balls for in-sample inference. For each sample element $(X_{i},Y_{i})\in\mathcal{L}_{n}$, begin by replacing the RF prediction of $X_{i}$ with its OOB prediction as the center of the ball. For each $(X_{j},Y_{j})\in\mathcal{L}_{n}\setminus\{(X_{i},Y_{i})\}$, compute the prediction based on the RF comprised of the trees in which both $(X_{i},Y_{i})$ and $(X_{j},Y_{j})$ are OOB, and use the predictions to obtain the ball radius. Although the effective number of trees is reduced, only one forest is trained, using the complete sample, which is an advantage over rank-one-out procedures for SC inference, especially for small sample sizes.
Our theoretical analysis assumes independence between prediction errors and predictors. Extensions to heteroscedastic settings that preserve asymptotic coverage guarantees are a natural direction for future work. Additionally, the scope of our work is currently limited to standard regression trees within RFs. An interesting extension would be to incorporate more sophisticated tree-based methods whose output goes beyond constant predictions by adapting, e.g., linear RFs (Raymaekers et al.,, 2024) to metric spaces through global Fréchet regression. Future work could also focus on deriving finite-sample guarantees for OOB balls.
@@ Appendix A Proofs of the main results
The second claim follows in a similar manner to Lemma $1$ in Zhang et al., (2020). We want to prove that $F_{R}\big(\widehat{R}_{[n,\gamma]}\big)\stackrel{{\scriptstyle\mathsf{P}}}{{\rightarrow}}\gamma$ as $n\to\infty$ for any $\gamma\in(0,1)$. We define
@@ C.2 Illustration of OOB balls on different metric spaces
Figure 12: Illustration of OOB balls in $\mathcal{S}_{2}^{+}$ for the example discussed in Section 3.3.3, with $d=2$ and $\alpha=0.10$. SPD matrices are represented via their eigen-decomposition: eigenvectors determine the orientation of the ellipse axes, and eigenvalues determine their lengths. The visualization is inspired by Figure 3.3 in Chacón and Duong, (2018).

### v2

@@ 3.3.4 Probability distributions on the real line
The same idea characterizes the Fréchet mean of $\mathrm{P}\in\mathcal{W}_{2}(\mathbb{R})$ as the unique measure $\mathrm{P}_{\oplus}\in\mathcal{W}_{2}(\mathbb{R})$ whose quantile function satisfies $F_{\mathrm{P}_{\oplus}}^{-1}=\mathsf{E}(F_{\mathrm{P}}^{-1})$ (see, e.g., Theorem 3.2.11 of Panaretos and Zemel, 2020), thus ensuring (c.2)(c.2.1).
The definition of regression models with response on $W_{2}(\mathbb{R})$ is relatively straightforward. For a given probability distribution $\mathrm{Y}\in\mathcal{W}_{2}(\mathbb{R})$, consider a slight modification of the regression model in Petersen and Müller (2019, Section 6.2):
@@
for some $\mu\in\mathcal{W}_{2}(\mathbb{R})$ and some finite real functions $f$ and $g$, under the restriction that $\sigma_{0}+g(x)>0$ for all $x$ in the support of $X$, so that $m(x)(\cdot)$ is a valid quantile function in $L^{2}[0,1]$. In order to generate the response, consider $F_{\mathrm{Y}}^{-1}(\cdot)=\gamma+f(X)+(\sigma+g(X))F_{\mu}^{-1}(\cdot)$, with $\gamma$ and $\sigma$ random variables independent of $X$ and following distributions of mean $\gamma_{0}$ and $\sigma_{0}$, respectively. See Serrano and García-Portugués (2025) for a detailed numerical example with this regression model. Thus, this general model satisfies (c.2)(c.2.2), and the conditional Fréchet mean is $m_{0}(x)(\cdot)$. The population OOB balls in this setting are formed by the distributions whose quantile functions are within a constant $L_{2}$-distance on $[0,1]$ equal to $R_{1-\alpha}$. If we assume, for simplicity, that $F_{\mu}^{-1}(\cdot)$ is the quantile of a distribution with mean zero and unit variance, then $R_{1-\alpha}$ is the $(1-\alpha)$-quantile of $((\gamma-\gamma_{0})^{2}+(\sigma-\sigma_{0})^{2})^{1/2}$.
@@ 5 Numerical experiments
In this section, we assess the performance of OOB balls through simulations with responses in the Euclidean space $\mathbb{R}^{q}$, on the unit sphere $\mathbb{S}^{2}$ and unit hyperboloid $\mathbb{H}^{2}$, and in the space of SPD matrices $\mathcal{S}_{2}^{+}$. In $\mathbb{R}^{q}$, OOB balls are compared with SC balls for metric data (Lugosi and Matabuena, 2024) in terms of coverage (Types I–IV), radius, and computational time.
@@
Each tree was trained with a size $n$ subsample generated with nonparametric bootstrap with replacement. We used $B=200$ trees, so that the number of trees in each OOB prediction is closer to the default parameter $B=100$ in scikit-learn (Pedregosa et al., 2011) and works related to RFs in metric spaces (Bulté and Sørensen, 2024; Qiu et al., 2024), without significantly increasing the computational cost of fitting and predicting. We investigated the effect of increasing $B$ on Type I coverage results, for OOB and SC balls, and found no significant improvement in coverage when the number of trees exceeded $B=200$.
The CART criterion was selected as impurity measure, and the split method used was the $2$-means algorithm. The minimal split size, which determines the minimal number of samples allowed in the leaf nodes, was tuned from $\{1,5,10\}$ using grid-search $5$-fold cross-validation with the MSE as error measure. Using the same cross-validation procedure, the number of random candidate variables to consider at each split was tuned from $\{1,2,\ldots,p\}$. This parameter is denoted by mtry in the pyfréchet package and in the most popular R programming language implementations (Liaw and Wiener, 2002; Wright and Ziegler, 2017). Three different values were considered for the significance level $\alpha$: $0.01,0.05$, and $0.10$.
@@ 5.2 Sphere
In the subsequent simulation scenarios, the focus is on the performance of OOB balls (radius and coverage types). We used RFWLCFR to train RFs using the implementation from Bulté and Sørensen (2024) in the pyfréchet package for Python. For the unit sphere $\mathbb{S}^{2}$, we considered the regression function $\boldsymbol{m}_{0}$ (great circle) given in (13), with $\boldsymbol{\mu}=(1/\sqrt{2},1/\sqrt{2})^{\top}$. The predictor variable $\Theta$ was generated as $\Theta=\operatorname{arctan2}(X_{2},X_{1})$, where $\boldsymbol{X}=(X_{1},X_{2})\sim\mathrm{vMF}((1,0)^{\top},1)$. For the response, we considered $\boldsymbol{Y}\mid\Theta=\theta\sim\mathrm{vMF}(\boldsymbol{m}_{0}(\theta),\kappa)$ for $\kappa=50$, $200$. The data generation process is illustrated in Figure 2 for different values of $\Theta$.
@@ 5.4 Symmetric positive definite matrices
We consider the space $\mathcal{S}_{q}^{+}$ of SPD $q\times q$ matrices, and study the impact of metric selection on the performance of OOB balls. We compared the coverage and radius of OOB balls in $\mathcal{S}_{2}^{+}$ using the three metrics. The regression model is constructed based on the idea of a convex interpolation between SPD matrices introduced in Section 3.3.3. In particular, we considered the interpolation function
@@ 7 Out-of-bag balls to model sunspot dynamics
Due to its plasma composition, the Sun’s equatorial regions rotate faster than the polar regions (differential rotation). This phenomenon distorts the magnetic field, stretching the Sun’s initial poloidal magnetic field into toroidal configurations, as plasma drags field lines along their rotational trajectories. Each solar cycle lasts around $11$ years and culminates when these twisted fields undergo a global polarity reversal (see, e.g., Babcock, 1961). The entanglement of field lines creates localized regions of intensified magnetic pressure that suppress convective heat transfer, manifesting as cooler, darker sunspots. Sunspots have a variable lifetime and serve as direct observational tracers of solar activity, which is linked to space weather events that impact spacecraft, astronauts, and Earth’s infrastructure.
We analyzed data based on a merged catalog combining the Greenwich Photoheliographic Results (GPR) and the Debrecen Photoheliographic Data (DPD) sunspot records (Baranyi et al., 2016; Győri et al., 2016). The GPR spans the years 1872–1976, while the DPD extends the record from 1974 onward, covering 14 solar cycles to date. The R package rotasym (García-Portugués et al., 2025) provides the data sets sunspots_births and sunspots_deaths, containing, respectively, the location on the Sun of the first and last record of each sunspot group (referred to as “birth” and “death” henceforth). We merged these data sets and excluded the first ($11$th) and last ($24$th) cycles due to the low reliability of their observations. Sunspots that were born and died on the same day were also excluded.
@@
For each solar cycle, we predicted death locations from birth locations using RFWLCFR, treating both response and predictors as points on the unit sphere $\mathbb{S}^{2}$ with geodesic distance $d_{\mathbb{S}^{2}}$. Although Qiu et al. (2024) consider only real predictors, their method can be applied straightforwardly when the predictors belong to any metric space. The data from each cycle were split into training ($75\%$) and test ($25\%$) sets, and the RFs were tuned with the same parameters as in Section 5, using the implementation in pyfréchet. Although the RF captured the latitude-dependent directional pattern (Figure 6(b)), the predicted longitudinal displacements are systematically smaller than observed in the training data. This shrinkage effect likely arises from the model’s averaging of opposing trajectories present at all latitudes, a consequence of the chaotic motion of sunspots.
@@
The predictors remained on $\mathbb{S}^{2}$ with the geodesic distance $d_{\mathbb{S}^{2}}$, while the responses were modeled on $S_{a,c}^{2}$, equipped with its intrinsic geodesic distance. Karney (2013) provides an algorithm to compute $d_{S_{a,c}^{2}}$ which can be made arbitrarily accurate and always converges. Although $d_{S_{a,c}^{2}}$ defines a Riemannian manifold structure on $S_{a,c}^{2}$, to simplify the implementation, we employed Fréchet medoids instead of Fréchet means in the CART criterion (4) and in the aggregation of results (i.e., we minimized (3) over the sample elements). On the sphere ($a=c=1$), medoids were also used for the sake of homogeneity in the comparison.
@@
Table 6: For each specification of $a$ (and $c=1$), the errors, areas, and coverages were calculated on the test points of cycles 21212323, and the reported values are weighted means across the three cycles, with weights proportional to the number of test points on each cycle. The areas correspond to OOB balls centered on the predicted value of the test points. The relative MSE difference $\Delta_{\mathrm{MSE}}$ and the relative mean area difference $\Delta_{\mathrm{area}}$ (both in $\%$) were calculated with respect to the balls on the sphere. The reported coverage corresponds to Type II (in %) for $\alpha=0.10$. The $p$-value of the one-sided paired $t$-test with alternative hypothesis $H_{1}:\mathrm{MSE}_{\mathbb{S}^{2}}<\mathrm{MSE}_{S_{a,1}}$ is $p_{\mathrm{MSE}}$. For $p_{\mathrm{area}}$, we considered a one-sided paired $t$-test with alternative $H_{1}:\mathrm{area}_{\mathbb{S}^{2}}>\mathrm{area}_{S_{a,1}}$, to test the equality of mean areas. False discovery rate correction Benjamini and Yekutieli (2001) is applied to the $p$-values at level $\alpha=0.01$.
@@ 8 Discussion
OOB prediction balls provide a principled approach to uncertainty quantification for RF predictions in metric data. The nature of a RF organically provides the OOB balls, with a main advantage over existing SC methods: the entire sample is used to train the RFs that generate the OOB balls, and to estimate the ball radius. Thus, OOB balls need half the sample size that standard SC methods require to achieve the same prediction and inference accuracy, as corroborated in simulations. Although OOB balls reduce the effective number of trees per observation by a factor of $e^{-1}$, this can be compensated by increasing the total number of trees.
The extension of the Euclidean prediction intervals (Zhang et al., 2020) to the metric setting introduces theoretical and computational challenges. From the theoretical viewpoint, the standard additive-error models are unavailable in the absence of algebraic structure, so they are adapted through conditions (c.2)(c.2.1) and (c.2)(c.2.2). The choice of metric space affects the performance of the RFs and the geometry of the OOB balls. In Section 7, we exploit this flexibility by constructing a data-driven distance. From the computational viewpoint, the additional cost does not stem from the construction of OOB balls itself, but from extending RFs to metric spaces (see Section 2.2). The estimation of Fréchet means is the main computational overhead, and depends strongly on the geometry of the metric space and on the availability of algorithms to compute them. Fréchet medoids provide a practical alternative, as evidenced in Section 7, which also represents, to the best of our knowledge, the first application of RFWLCFR with metric predictors.
OOB balls also offer a natural path toward in-sample inference. Suppose one wants to analyze the confidence regions at the same points where the RF was trained. Rank-one-out procedures (Lei et al., 2018) reduce the cost of in-sample SC inference, but still inherit the inefficiency of SC prediction due to halving the sample size, and the computational cost is doubled since one RF is trained on each half sample. A possible alternative is to apply OOB balls for in-sample inference. For each sample element $(X_{i},Y_{i})\in\mathcal{L}_{n}$, begin by replacing the RF prediction of $X_{i}$ with its OOB prediction as the center of the ball. For each $(X_{j},Y_{j})\in\mathcal{L}_{n}\setminus\{(X_{i},Y_{i})\}$, compute the prediction based on the RF comprised of the trees in which both $(X_{i},Y_{i})$ and $(X_{j},Y_{j})$ are OOB, and use the predictions to obtain the ball radius. Although the effective number of trees is reduced, only one forest is trained on the complete sample, which is an advantage over rank-one-out procedures for SC inference, especially for small sample sizes.
Our theoretical analysis assumes independence between prediction errors and predictors. Extensions to heteroscedastic settings that preserve asymptotic coverage guarantees are a natural direction for future work. Additionally, the present work focuses on standard regression trees within RFs. Adapting richer tree-based models to metric spaces, such as linear RFs (Raymaekers et al., 2024) via global Fréchet regression, is another promising avenue. Future work could also focus on deriving finite-sample guarantees for OOB balls.
@@ Appendix A Proofs of the main results
The second claim follows in a similar manner to Lemma $1$ in Zhang et al. (2020). We want to prove that $F_{R}\big(\widehat{R}_{[n,\gamma]}\big)\stackrel{{\scriptstyle\mathsf{P}}}{{\rightarrow}}\gamma$ as $n\to\infty$ for any $\gamma\in(0,1)$. We define
@@ C.2 Illustration of OOB balls on different metric spaces
Figure 12: Illustration of OOB balls in $\mathcal{S}_{2}^{+}$ for the example discussed in Section 3.3.3, with $d=2$ and $\alpha=0.10$. SPD matrices are represented via their eigen-decomposition: eigenvectors determine the orientation of the ellipse axes, and eigenvalues determine their lengths. The visualization is inspired by Figure 3.3 in Chacón and Duong (2018).